OpenStudy (aaronandyson):

A refrigerator converts 100g of water at 20 C to ice at - 10 C in 73.5 mins. Calc the average rate of heat extraction in watt. specify heat capacity of water is 4.2 J g^-1 K^-1.specific latent heat capacity of ice is 336 J g ^-1

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

here we have to compute the amount of heat which has to be extracted, first

OpenStudy (aaronandyson):

How to do so..??

OpenStudy (michele_laino):

such amount of heat is: \[\large Q = Lm + cm\Delta \theta = \left( {336 \cdot 100} \right) + \left( {4.2 \cdot 100 \cdot 30} \right) = ...?\]

OpenStudy (aaronandyson):

46200

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

so, the requested average rate \(r\), is given by the subsequent computation: \[\Large r = \frac{Q}{{\Delta t}} = \frac{{46200}}{{\left( {73.5 \cdot 60} \right)}} = ...Watt\]

OpenStudy (aaronandyson):

10.47

OpenStudy (michele_laino):

since we need to measure time in seconds, here is why I wrote \(73.5 \cdot 60\)

OpenStudy (aaronandyson):

10.47 Watts..?

OpenStudy (michele_laino):

correct!

OpenStudy (aaronandyson):

73.5*60 = 4410

OpenStudy (michele_laino):

we can write: \(r=10.5 \; watts\)